Project

General

Profile

Commissioning and Ops Scheduling meeting : 2/2/2016

To be held in CR10 in the ATLAS Building @ 15:30 pm

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

Agenda:

  • Review of tasks and activities identified at last meeting
    • Action Item : John will look at the effect of air on the MC truth scattering angle
    • Action Item : Pierrick/Mark - ensure that the pressure and material in the absorber volumes are recorded
      • Action Item : DC to work with Pierrick to ensure this
    • Action Item : Steve to talk to Alan Stevens about stopping 4-hourly monitoring but leaving the alarm in place
    • Action Item : Paolo, Mike and Vicky will investigate the Linde fridge PC instability issue
    • Action Item : Paolo/Pierrick will think of a way to monitor the Linde fridge PC
    • Action Item : Colin, Trevor, Steve, Pierrick to have a meeting to review the Daresbury work schedule in light of running time
      and the currently known state of the SS work. Meeting to happen on morning of Thursday 28th.
  • Current schedule
    Feb 1 - 5 : Daresbury Electrical work ( 10 A test )
    Feb 2 : DS coldbox cool down
    Feb 4 : DS magnet cool down
  • Updates and future schedule
    • DS Cooldown schedule
    • Realignment work and schedule
    • Cooling water discussion
    • Controls and Monitoring : Network reconfiguration
    • Tracker patch panel leak work
    • Electrical work schedule
    • Data-taking schedule
      • User cycle begins Feb 16th
      • Start-up activities : Feb 16 - 21
      • Data-taking starts : Feb 23

Minutes

Present: Steve B, Chris R, Paul H, Andy N, Vicky B, Mike C, Ryan B, Paul K, Paolo F, Josef F, Henry N, Mark T, Melissa U, Victoria B, John N

*Action Items
  • Action Item : John will look at the effect of air on the MC truth scattering angle
    • John presented plots of MC scattering angle in air and vacuum. They contain some features that we don't understand. John will continue looking into these.
  • Action Item : Pierrick/Mark - ensure that the pressure and material in the absorber volumes are recorded
    • Action Item : DC to work with Pierrick to ensure this
    • Action Item stands
  • Action Item : Steve to talk to Alan Stevens about stopping 4-hourly monitoring but leaving the alarm in place
    • Email sent but no reply. Steve will try again
  • Action Item : Paolo, Mike and Vicky will investigate the Linde fridge PC instability issue
    • The IOC needed to be restarted. This fixed the problem
  • Action Item : Paolo/Pierrick will think of a way to monitor the Linde fridge PC
    • Paolo has added a monitoring process to Nagios to ensure that the IOC on this PC is always alive
  • Action Item : Colin, Trevor, Steve, Pierrick to have a meeting to review the Daresbury work schedule in light of running time
    and the currently known state of the SS work. Meeting to happen on morning of Thursday 28th.
    • Action Item postponed. However we do need to do this in the near future. Colin and Trevor spoke informally about priorities and Trevor looked through the schedule. The FC requires about 2 weeks of work. Once it's online some work becomes near-impossible. Trevor is focussing on the FC now. DL are also helping out on the specs of the dump resistors. Work being done on the SS is conditional on the results Steve gets from 10A test. Pierrick reported that the 10A tests are finishing up now. There is no real conclusion on this test. Did find that a small resistance in parallel to the magnet caused overshoots and undershoots so not properly damped. If issues can't be solved we may have to change where the resistors go. Next week most of the PS modification hardware is being sent to RAL so most of the work is done before March is done.
    • Issues of chain of authority during "mixed working" i.e. work in rack room at the same time as data-taking needs to be addressed.
    • We all agreed that the DC is in overall charge, even during data-taking. However, we need think about the key dance and what needs to
      be done to ensure clarity in the line of authority.
    • Action Item : Steve/DC/Andy to think about organisation during mixed work sessions
  • Current activities
  • DS Cooldown schedule
    • DS will start cooldown on Feb 3
  • Realignment work and schedule
    • Drawing for modified bellows is ready to go out, but the bellows probably won't be ready for Feb 23 when data-taking begins.
    • All decided to do as much work as we can on the realignment but not move the solenoids now.
    • The actual realignment work will take place between the next and next-2-next user cycles once everything is ready and procedure is understood.
  • Tracker patch panel leak work
    • We need to know what the plan for work on the He leak in the tracker is by next week as this work needs to be done before other items.
    • Tracker group to be asked to think about their plan and report at next Ops meeting
    • Action Item Steve/Colin to talk to the tracker group and request action plan
  • Realignment of PID detectors
    • This has been done and alignment guides installed on the floor.
    • EMR needs to go in position, but this can't happen until we understand the plan for fixing the tracker He leaks
    • There is a question of how we run if the SC permit is required but the PRY is open (e.g. at low current).
    • Action Item Andy/Colin/DC to think about this procedure
  • Cooling water discussion
    • This discussion has been parked for the time being. Will need to return to it after the user cycle finishes.
  • Controls and Monitoring : Network reconfiguration
    • All IOCs configured with static IPs. Now need to test each IOC. Network lockout code is being written. Will be done by the next meeting.
  • Electrical work schedule
    • Action Item Schedule a meeting between Colin/Trevor/Myself/Pierrick.
  • Data-taking schedule
    • Was explained in order to Focus minds

air_drift_proj_x_mc.pdf (13.5 KB) Nugent, John , 02 February 2016 15:26

vac_drift_proj_x_mc.pdf (13.6 KB) Nugent, John , 02 February 2016 15:26